OpenStudy (anonymous):

find the vertex:

OpenStudy (anonymous):

\[x=2y^2-4y=1\]

OpenStudy (anonymous):

x=2y^2-4y+1*

OpenStudy (dumbcow):

2 ways of doing this 1) use completing the square to get it in vertex form 2) use shortcut formula, y = -b/2a where a,b,c refer to the coefficients

OpenStudy (anonymous):

I wasn't sure if I was allowed to work it the same way as I solved for y.

OpenStudy (dumbcow):

hmm you don't need to solve for y, just get the vertex a = 2 b=-4 c=1 y = 4/4 = 1 plug that in to get x_coordinate

OpenStudy (anonymous):

So it should work out to V(1,-1)?

OpenStudy (dumbcow):

other way around, (x,y) --> (-1,1) :)

OpenStudy (anonymous):

Oh, good point Thank you!